Add in the indicated base
221(sub)4 + 131 (sub)4

*sub indicates the smaller numbers after the 1 in both numbers.*

Answer by solver91311(24713)



One and one is two, two is less than 4, so

   221
 + 131
------
     2

2 and 3 is 5, but 5 is one 4 plus one 1, bring down 1, carry 1

   1
   221
 + 131
------
    12

1 and 2 and 1 is 4, bring down 0, carry 1

  11
   221
+  131
------
   012

1 and 0 is 1, bring down 1

  11
   221
+  131
------
  1012


Check:

1 + 2 * 4 + 2 * 16 = 1 + 8 + 32 = 41, so 221 base 4 is 41 base 10
1 + 3 * 4 + 16 = 1 + 12 + 16 = 29, so 131 base 4 is 29 base 10.
And 41 base 10 plus 29 base 10 is 70 base 10

If the problem above was done correctly, then 1012 base 4 should be 70 base 10.
2 + 4 + 0 + 64 = 70.

Checks.


John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism